OpenStudy (anonymous):

how to find r when h=7, k= -3 touches the x axis????

OpenStudy (anonymous):

\[(x-h)^{2} + (y-k)^2 = r^2\]

OpenStudy (anonymous):

(x−h)2+(y−k)2=r2